OpenStudy (anonymous):

Simplify the expression, if possible (Rewarding medals) log 10^5 log 10 log 10^-4 log 10^-1 log 1 log 3√100 log 1/100 log 1/10 log 10^4.7 log 10^2x+4

OpenStudy (anonymous):

I'm not going to answer all these, but I'll show you a few of the log rules you'll be using in simplifying them. First off: \[\log(x^a)=alog(x)\]In other words, when you have a base raised to a power inside a log, the exponent can drop down in front. So, in your first example: log (10^5)=5log(10)

OpenStudy (anonymous):

The other rules you'll want are the addition and subtraction of logs: \[\log(ab) = \log(a) + \log(b)\]If we have two quantities multiplied together inside a log, we can split it into two logs, one for each quantity, added together. \[\log(\frac{ a }{ b }=\log(a) - \log(b)\]If we have division, then we do the same thing, but we subtract the logs.

OpenStudy (anonymous):

I have worked them out I basically want to check my answers... log 10^5= 5 log 10=1 log 10^-4=0.0001 log 10^-1= 0.1 log 1= 0 log 3√100=3

OpenStudy (anonymous):

The first two are correct, the next two are not. It looks like you simplified the 10^-4 and wrote it in decimal form, then just forgot about the log. You could use the first rule I showed you, since you have an exponent inside. So put the -4 to the front: -4log10=-4(1)=-4